# Agent Contracts: A Formal Framework for Resource-Bounded Autonomous AI Systems (Full)

### v1

@@ Abstract
[…]
@@ 4.3 Contract Lifecycle
Contracts progress through distinct states following the transition pattern DRAFTED $\rightarrow$ ACTIVE $\rightarrow$ {FULFILLED, VIOLATED, EXPIRED, TERMINATED}. A contract begins in the DRAFTED state, where its parameters are specified but execution has not begun. Activation transitions the contract to ACTIVE, at which point resources are reserved and monitoring begins. From ACTIVE, the contract reaches exactly one of four terminal states.
@@ 7.2 Enforcement Capabilities
Contracts provide particularly high value in five scenarios. First, retry loops benefit from contracts that prevent runaway costs when tasks fail repeatedly. Second, validation cycles benefit from contracts that limit iterative refinement (e.g., code generation $\rightarrow$ test $\rightarrow$ fix loops). Third, multi-agent workflows benefit from contracts that stop downstream agents when upstream exceeds budget. Fourth, tool-heavy agents benefit from contracts that control cumulative cost of web searches and API calls. Fifth, long-running sessions benefit from contracts that enforce session-level budgets across many interactions.
@@ 8.2 Runaway Prevention in Iterative Workflows
[…]
@@
[…]
@@
Token Usage 34,606 3,461 -90% 0.0007***
[…]
Iterations 3.00 1.71 -43% <0.0001***
LLM Calls 9.0 4.5 -50% <0.0001***
Success Rate 60.0% 52.9% -7.1pp 0.13 (NS)
@@ 8.3 Conservation Laws in Multi-Agent Coordination
We evaluate conservation laws and contract delegation (Section 6) using a three-agent research pipeline: Researcher $\rightarrow$ Analyzer $\rightarrow$ Reporter. The orchestrator delegates sub-contracts to each worker via DelegatingAdkAgent, enforcing both the conservation invariant $\sum b_{i}\leq B$ and per-tool limits at allocation time.
@@
[…]
@@ 8.4 Quality-Resource Tradeoffs via Contract Modes
Design. Fifty medium-difficulty logic puzzles are evaluated under three contract modes: URGENT (no extended reasoning, 30s timeout), ECONOMICAL (low reasoning effort, 60s timeout), and BALANCED (medium reasoning effort, 90s timeout). Success requires exact numeric answer match.
@@
Framework Component Experiment Validation
Budget awareness All experiments Total budget in prompts; dynamic status updates (Code Review)
Resource constraints $R$ Code Review Tokens, iterations, LLM calls
Agent delegation Research Pipeline Sub-contracts, conservation laws, per-tool limits
Success criteria $\Phi$ Crisis Comm. Quality thresholds
Contract modes Strategy Modes URGENT/ECONOMICAL/BALANCED
@@ 9 Conclusion
[…]

### v2

@@ Abstract
The Contract Net Protocol (1980) introduced coordination through contracts in multi-agent systems. Modern agent protocols standardize connectivity and interoperability, yet none provide formal resource governance—normative mechanisms to bound how much agents may consume or how long they may operate. We introduce Agent Contracts, a formal framework that extends the contract metaphor from task allocation to resource-bounded execution. An Agent Contract $C=(I,O,S,R,T,\Phi,\Psi)$ unifies input/output specifications, multi-dimensional resource constraints, temporal boundaries, and success criteria into a coherent governance mechanism with explicit lifecycle semantics. For multi-agent coordination, we establish conservation laws ensuring delegated budgets respect parent constraints, enabling hierarchical coordination through contract delegation. Empirical validation across four experiments demonstrates 90% token reduction with 525× lower variance in iterative workflows, zero conservation violations in multi-agent delegation, and measurable quality-resource tradeoffs through contract modes. Agent Contracts provide formal foundations for predictable, auditable, and resource-bounded autonomous AI deployment.
@@ 4.3 Contract Lifecycle
Contracts progress through distinct states following the transition pattern DRAFTED → ACTIVE → {FULFILLED, VIOLATED, EXPIRED, TERMINATED}. A contract begins in the DRAFTED state, where its parameters are specified but execution has not begun. Activation transitions the contract to ACTIVE, at which point resources are reserved and monitoring begins. From ACTIVE, the contract reaches exactly one of four terminal states.
@@ 7.2 Enforcement Capabilities
Contracts provide particularly high value in five scenarios. First, retry loops benefit from contracts that prevent runaway costs when tasks fail repeatedly. Second, validation cycles benefit from contracts that limit iterative refinement (e.g., code generation → test → fix loops). Third, multi-agent workflows benefit from contracts that stop downstream agents when upstream exceeds budget. Fourth, tool-heavy agents benefit from contracts that control cumulative cost of web searches and API calls. Fifth, long-running sessions benefit from contracts that enforce session-level budgets across many interactions.
@@ 8.2 Runaway Prevention in Iterative Workflows
The “\$47K problem” from the Introduction illustrates the danger of unbounded agent loops. We evaluate iteration governance using a Coder↔Reviewer pipeline where agents iteratively refine code solutions. The Coder writes Python code; the Reviewer executes it against test cases using a test_code tool and either approves or requests revision. The experiment uses 70 problems from LiveCodeBench [27] (released post-February 2025, after model cutoff): 31 easy and 39 medium difficulty.
@@
Results. CONTRACTED execution achieves 90% token reduction compared to UNCONTRACTED ($p=0.0007$, paired $t$-test), with 525× lower variance (5.29B vs 10.1M)—directly addressing the “\$47K problem.” The success rate difference of 7.1 percentage points (52.9% vs 60.0%) is not statistically significant ($p=0.13$). Governance value increases with task complexity: medium-difficulty problems show 92% token savings versus 76% for easy.
@@
Token Usage 34,606 3,461 −90% 0.0007***
Variance 5.29B 10.1M 525× lower
Iterations 3.00 1.71 −43% <0.0001***
LLM Calls 9.0 4.5 −50% <0.0001***
Success Rate 60.0% 52.9% −7.1pp 0.13 (NS)
@@ 8.3 Conservation Laws in Multi-Agent Coordination
We evaluate conservation laws and contract delegation (Section 6) using a three-agent research pipeline: Researcher → Analyzer → Reporter. The orchestrator delegates sub-contracts to each worker via DelegatingAdkAgent, enforcing both the conservation invariant $\sum b_{i}\leq B$ and per-tool limits at allocation time.
@@
Results. CONTRACTED execution achieves zero conservation violations across all 50 trials, demonstrating that conservation laws ($\sum b_{i}\leq B$) are fully enforceable. In one trial, the enforcement mechanism successfully detected and halted a runaway agent that exceeded its 40K token budget (56K consumed), providing evidence that runtime enforcement works as designed. Quality variance is 26.7× lower than UNCONTRACTED ($\sigma$: 1.75 vs 9.07), though partly driven by one catastrophic failure in the UNCONTRACTED condition. Excluding this outlier, CONTRACTED still shows 1.4× lower variance (88.5% Bayesian probability). The key insight is not mean quality improvement but variance reduction—contracts eliminate catastrophic failures where agents consume resources without producing useful output.
@@ 8.4 Quality-Resource Tradeoffs via Contract Modes
Design. Fifty medium-difficulty logic puzzles are evaluated under three contract modes: URGENT (no extended reasoning, 8s timeout), ECONOMICAL (low reasoning effort, 10s timeout), and BALANCED (medium reasoning effort, 30s timeout). Success requires exact numeric answer match.
@@
Across all four experiments, the consistent finding is not cost optimization but governance: contracts transform unpredictable agent behavior into bounded, auditable operations. Code Review demonstrates multi-dimensional resource enforcement (90% token reduction, 525× variance reduction). Research Pipeline validates conservation laws and runtime enforcement (zero conservation violations; one runaway agent detected and halted). Crisis Communication shows that quality thresholds prevent failures, not just reduce costs. Strategy Modes confirms that contract configurations operationalize satisficing tradeoffs (70%→86% success).
Framework Component Experiment Validation Specification
Budget awareness All Prompts; dynamic status "Budget: {used}/{total}"
Resource constraints $R$ Code Review Tokens, iterations, calls $r_{tok}$=50K, $r_{iter}$=3, $r_{llm}$=6
@@
Success criteria $\Phi$ Crisis Comm. Quality thresholds $Q_{min}=0.80$
Contract modes Strategy URGENT/ECON./BALANCED $\tau$: 8s / 10s / 30s
@@ 9 Conclusion
Empirical validation demonstrates that the formal framework translates into practical governance: resource constraints enable 90% token reduction with 525× lower variance; conservation laws achieve 100% compliance in multi-agent delegation; and contract modes operationalize satisficing tradeoffs (70%→86% success).